Air enters the compressor of a gas turbine at 100 kPa and 25°C. For a pressure ratio of 5 and a maximum temperature of 850°C determine the back work ratio and the thermal efficiency using the Brayton cycle.



1
Expert's answer
2021-09-19T00:31:08-0400


P1=100kPa

T1=25°C=298K

r=5

T3=850°C=1123K

at T1 and P1,

h1= "295.17+(\frac{298-295}{300-295})*(300.9-295.17)"

=298.608kJ/kg

(pr)1="1.3068+(\frac{298-295}{300-295})*(1.386-1.3068)"

=1.35432

(pr)2="(\frac{P_2}{P_1})*(p_r)_1"

=5*1.35432

=6.7716

At (pr)2 by interpolation,

h2="472.24+(\frac{6.7716-6.742}{7.268-6.742})*(482.49-472.24)"

=472.8186kJ/kg

At T3,

h3="1184.28+(\frac{1123-1120}{1140-1120})*(1207.57-1184.28)"

=1187.77kJ/kg

(pr)3="179.7+(\frac{1123-1120}{1140-1120})*(193.1-179.7)"

=181.71

(pr)4="(\frac{P_4}{P_3})*(p_r)_3"

="\frac{1}{5}*181.71"

=36.342

At (pr)4 by interpolation,

h4="756.44+(\frac{36.342-35.5}{37.35-35.5})*(767.29-756.44)"

=767.38kJ/kg

Back work ratio ="\frac{W_c}{W_t}=\frac{h_2-h_1}{h_3-h_4}"

="\frac{472.8168-298.608}{1187.77-767.38}"

=0.4144

bwr=41.44%

Efficiency ="\frac{(1187.77-767.38)-(472.8168-298.608)}{1187.77-472.8168}"

=0.344

=34.4%